An industrial plant produces 150 cars in 6 weeks. If the plant operates 5 days a week, how many cars does it produce each day?

Thank you!
1
Expert's answer
2019-09-13T10:32:02-0400

"\frac{150 \text{ cars}}{6 \text{ week} \cdot 5 \frac{\text{day}}{\text{week}}} = \frac{150 \text{ cars}}{30 \text{ day}} = 5 \frac{\text{cars}}{\text{day}}."